# Association and Independence Test for Random Objects

### v2

@@ 5 Simulations
We report here finite sample comparisons of the performance of the proposed profile approach with various alternative methods when $X$ and $Y$ are located in pertinent metric spaces. These include the Euclidean space $\mathbb{R}^{p}$, the sphere $\mathbb{S}^{2}$ and the 2-Wasserstein space of one-dimensional distributions $\mathbb{W}^{2}$ with the Wasserstein metric, where the latter is
$d_{W}(x,y)^{2}=\int_{0}^{1}\{Q_{x}(u)-Q_{y}(u)\}^{2}\,{\rm d}u$ (16)
and $Q_{x},\,Q_{y}$ are the quantile functions associated with the distributions $x$ and $y$.
@@
For each setting, both the proposed and comparison methods were implemented using the permutation test described in Section 3.2 and empirical power curves at significance level 0.05 are obtained from 500 Monte Carlo runs. We consider the following settings, where in all settings the dependence is quantified by the parameter $\rho$, where $\rho=0$ corresponds to independence and the level of dependence (association) increases with $\rho$. We obtained empirical power curves that illustrate how the power of the profile independence test increases as the argument $\rho$ increases, where we consider $0\leq\rho\leq 1$
  • $\mathbb{R}^{1}$-Cir: Denoting by $\|v\|_{\zeta}$ the $\zeta$-norm for any Euclidean vector $v$. For any $1\leq\zeta\leq\infty$ we consider pairs $(X,Y)$ of one-dimensional random variables $X$ and $Y$ that satisfy the condition $\|(X,Y)\|_{2/\rho}\leq 1$, where $X$ and $Y$ are independently drawn from $\text{Unif}(-1,1)$. When $\rho=0$, $(X,Y)$ is uniformly distributed over the square $[-1,1]\times[-1,1]$. When $\rho=1$, $(X,Y)$ follows a uniform distribution within the unit circle $\{(x,y)\mid x^{2}+y^{2}\leq 1\}$, with domains for the uniform distribution located in between the square and the circle for $10<\rho<$ 1 .
  • $\mathbb{R}^{1}$-Bimodal: The pair $(X,Y)$ of one-dimensional random variables satisfies $Y=\epsilon|X|^{\rho}$, where $X$ and $\epsilon$ are independent and follow mixture distributions $0.2\mathcal{N}(-3,1)+0.8\mathcal{N}(3,1)$.
@@
  • $\mathbb{R}^{p}$-Log: $\bm{X}=(X_{1},\ldots,X_{p})^{\top}$ and $\bm{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ are related only through their first elements, specifically $Y_{1}=\rho\log(4X_{1}^{2})+\epsilon$. The variables $Y_{2},\ldots,Y_{p}$, $X_{1},X_{2},\ldots,X_{p}$, and $\epsilon$ are all distributed as $\mathcal{N}(0,1)$ and are independent.
  • $\mathbb{R}^{p}$-Pow: $\bm{X}=(X_{1},\ldots,X_{p})^{\top}$ and $\bm{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ are related only through the first element, where $Y_{1}=\epsilon\times(|X_{1}|^{\rho})$. The variables $Y_{2},\ldots,Y_{p}$ and $\epsilon$ are distributed as $\mathcal{N}(0,1)$, while $X_{1},X_{2},\ldots,X_{p}$ are distributed as $\mathrm{Unif}(-1,1)$. Additionally, $Y_{2},\ldots,Y_{p}$, $X_{1},X_{2},\ldots,X_{p}$, and $\epsilon$ are independent.
  • $\mathbb{S}^{2}$-Lin: $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by independent standard normal random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$:
    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
    $Y_{1}=\frac{\rho\epsilon_{1}+(1-\rho)\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|},$
    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(\rho\epsilon_{1}+(1-\rho)\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
  • $\mathbb{S}^{2}$-Log: $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by independent standard normal random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$:
    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
    $Y_{1}=\frac{\rho\log(4\epsilon_{1}^{2})+\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|},$
    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(\rho\log(4\epsilon_{1}^{2})+\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
  • $\mathbb{S}^{2}$-Pow: $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by independent standard normal random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$:
    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
    $Y_{1}=\frac{|\epsilon_{1}|^{\rho}+\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|},$
    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(|\epsilon_{1}|^{\rho}+\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
  • $\mathbb{S}^{2}$-Pow-Bimodal: Here $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by the random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$, where $\epsilon_{i}$ are random samples from a mixture of Gaussians $0.2\mathcal{N}(-1,1)+0.8\mathcal{N}(1,1)$:
    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
    $Y_{1}=\frac{|\epsilon_{1}|^{\rho}+\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|}$
    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(|\epsilon_{1}|^{\rho}+\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
@@
  • $\mathbb{W}^{2}$-Mean: $X$ and $Y$ are random elements in the space of distributions $\mathbb{W}^{2}$ with the Wasserstein metric, specifically $X=\mathcal{N}(\mu_{X},1)$ and $Y=\mathcal{N}(\mu_{Y},1)$, where $\mu_{X}\sim\mathrm{Unif}(-1,1)$, $\mu_{Y}=\epsilon\times|\mu_{X}|^{\rho}$, and $\epsilon\sim\mathcal{N}(0,1)$ and is independent of $\mu_{X}$.
  • $\mathbb{W}^{2}$-Var: $X$ and $Y$ are random elements in the space of distributions $\mathbb{W}^{2}$ with the Wasserstein metric, specifically $X=\mathcal{N}(0,\sigma_{X}^{2})$ and $Y=\mathcal{N}(0,\sigma_{Y}^{2})$, where $\sigma_{X}^{2}\sim\mathrm{Unif}(0,1)$ and $\sigma_{Y}=\rho\sin(\sigma_{X}\pi)+\epsilon$, and where $\epsilon\sim\mathrm{Unif}(0,2)$ and is independent of $\sigma_{X}$.
For cases where $X$ and $Y$ are in $\mathbb{R}^{p}$ for $p\geq 1$, we compare the proposed profile independence test with the energy test (Székely et al., 2007), the ball divergence test (Pan et al., 2020) and Chatterjee’s rank test (Chatterjee, 2021). All methods are implemented using R packages: the energy test (Energy) via dcov.test() from the energy package, the ball divergence test (Ball) via bcov.test() through the Ball package (Zhu et al., 2021) and Chatterjee’s rank test (Chatterjee) via xicor() through the XICOR package. To compare different approaches, we focus on the power curves as function of $\rho$ for the simulation settings described above. Figure 1 indicates that the proposed method outperforms all comparable methods, particularly in the $\mathbb{R}^{1}$-Bimodal setting, where the energy test and Chatterjee’s rank test fail. Additionally, in this more complex setting, the proposed test exhibits a faster increase in power as the sample size increases.
Figure 2 shows that the proposed profile independence test (blue) continues to consistently outperform the energy test (black) and the ball divergence test (red) across all simulation settings where $X$ and $Y$ reside in multivariate Euclidean spaces or metric spaces. The simulations thus provide evidence that the profile independence test effectively adapts to complex dependency structures where other methods fail. For example, when the relationship between $X$ and $Y$ is linear, a scenario where also classical methods perform well, the energy test performs comparably to the profile independence test when the relationship between $X$ and $Y$ is linear. However, in more complex cases, such as log- or power-based dependencies, and especially in settings $\mathbb{S}^{2}$-Log and $\mathbb{S}^{2}$-Pow, the energy test performs poorly, whereas the profile independence test maintains strong detection power. These results provide evidence for the robustness and efficiency of the proposed test in detecting dependencies across diverse scenarios.
Figure 1: Power curves comparing the proposed profile independence test (blue), the ball statistic test (red), the Chatterjee est (yellow) and the energy test (black), where power is a function of the increasing dependency parameter $\rho$, for two simulation settings: $\mathbb{R}^{1}$-Cir (first row) and $\mathbb{R}^{1}$-Bimodal (second row) The dashed grey line indicates the target significance level of $0.05$. Three different sample sizes are considered: $n=100$ (left column), $n=200$ (middle column), and $n=400$ (right column).
@@
Figure 2: Power curves comparing the proposed profile independence test (blue), the ball statistic test (red), the Chatterjee test (yellow) and the energy test (black), where power is a function of the increasing dependency parameter $\rho$, for nine simulation settings: $\mathbb{R}^{5}$-Log, $\mathbb{R}^{10}$-Log, and $\mathbb{R}^{10}$-Pow (left, middle, right panel in first row); $\mathbb{S}^{2}$-Lin, $\mathbb{S}^{2}$-Log, and $\mathbb{S}^{2}$-Pow (left, middle, right panel in second row); $\mathbb{S}^{2}$-Pow-Bimodal, $\mathbb{W}^{2}$-Mean, and $\mathbb{W}^{2}$-Var (left, middle, right panel in third row). The dashed grey line shows the target significance level of $0.05$, and the sample size is set to $n=400$.

### v3

@@ 5 Simulations
We present finite-sample comparisons of the proposed test against several alternative methods when $X$ and $Y$ are located in different metric spaces. These include the Euclidean space $\mathbb{R}^{p}$, the sphere $\mathbb{S}^{p}$, the space of $d$-dimensional symmetric positive definite matrices $\mathcal{S}_{p}^{++}$, and the 2-Wasserstein space of one-dimensional distributions. For the Euclidean space $\mathbb{R}^{p}$ and the sphere $\mathbb{S}^{p}$, we use the usual $\mathcal{L}^{2}$ distance and the geodesic metric, respectively. For the $\mathcal{S}_{p}^{++}$ space, we use the affine-invariant Riemannian distance $d(A,B)=\|\mathfrak{log}(A^{-1}B)\|_{F}^{2}$, where $\mathfrak{log}$ and $\mathfrak{exp}$ are the logarithm and exponential functions for matrices. Moreover, for $A,B\in\mathcal{S}_{p}^{++}$, the affine-invariant geodesic interpolation is defined by
$A\#_{\rho}B=A^{1/2}(A^{-1/2}BA^{-1/2})^{\rho}A^{1/2}.$
For the Wasserstein space $\mathbb{W}^{2}$, we use the 2-Wasserstein metric, which is defined as
$d_{W}(x,y)^{2}=\int_{0}^{1}\{Q_{x}(u)-Q_{y}(u)\}^{2}\,{\rm d}u.$ (16)
@@
Here $Q_{x},\,Q_{y}$ are the quantile functions associated with the distributions $x$ and $y$. For each setting, both the proposed and comparison methods were implemented using the permutation-based test and empirical power curves at significance level 0.05 are obtained from 400 Monte Carlo runs.
For the unconditional testing problem in Section 3, we consider the following settings, where in all settings the dependence is quantified by the parameter $\rho$, where $\rho=0$ corresponds to independence and the level of dependence (association) increases with $\rho$. We obtained empirical power curves that illustrate how the power of the profile independence test increases as the argument $\rho$ increases, where we consider $0\leq\rho\leq 1.$
@@
  • $\mathbb{R}^{p}$-Lin: $\bm{X}=(X_{1},\ldots,X_{p})^{\top}$ and $\bm{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ are related only through their first $J_{p}=\max\{1,\lfloor 0.1p\rfloor\}$ components, where each $Y_{j}$ for $j\leq J_{p}$ is generated as $Y_{j}=\rho X_{j}+\varepsilon_{j},$ with $\varepsilon_{j}\sim\mathcal{N}(0,1)$ independent of all other variables. The remaining components $Y_{J_{p}+1},\ldots,Y_{p}$, and $X_{1},\ldots,X_{p}$, are independently distributed as $\mathcal{N}(0,1)$.
  • $\mathbb{R}^{p}$-Log: $\bm{X}=(X_{1},\ldots,X_{p})^{\top}$ and $\bm{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ are generated as in the $\mathbb{R}^{p}$-Lin setting, except that $Y_{j}=\rho\log(4X_{j}^{2})+\varepsilon_{j}$, where $\varepsilon_{j}\sim\mathcal{N}(0,1)$ is independent of all other variables.
  • $\mathbb{R}^{1}$-Cir: Denoting by $\|v\|_{\zeta}$ the $\zeta$-norm for any Euclidean vector $v$. For any $1\leq\zeta\leq\infty$ we consider pairs $(X,Y)$ of one-dimensional random variables $X$ and $Y$ that satisfy the condition $\|(X,Y)\|_{2/\rho}\leq 1$, where $X$ and $Y$ are independently drawn from $\text{Unif}(-1,1)$. When $\rho=0$, $(X,Y)$ is uniformly distributed over the square $1,1]\times[-1,1]$. When $\rho=1$, $(X,Y)$ follows a uniform distribution within the unit circle $\{(x,y)\mid x^{2}+y^{2}\leq 1\}$, with domains for the uniform distribution located in between the square and the circle for $0<\rho<1.$
  • $\mathcal{S}_{2}^{++}$-Interpolation: $X$ and $Y$ are $2\times 2$ symmetric positive definite matrices. $X$ is generated as $X=\mathfrak{exp}((S+S^{\top})/2)$, where $S$ is a $2\times 2$ matrix with independent entries drawn from $\mathcal{N}(0,0.6^{2})$. Let $Y_{(0)}$ denote the independent baseline, which is generated independently but follows the same distribution as $X$. The dependent target $Y_{(1)}$ incorporates geodesic noise with $P=CXC^{\top}$, and is constructed as
    $Y_{(1)}=P^{1/2}\mathfrak{exp}(0.1(E+E^{\top})/2)P^{1/2},$
    where $C={\rm diag}(2,0.5)$ and $E$ is a $2\times 2$ matrix with independent standard normal entries. The final matrix $Y$ is obtained via geodesic interpolation between the independent baseline $Y_{(0)}$ and the dependent target $Y_{(1)}$, defined as
    $Y=Y_{(0)}\#_{\rho/5}Y_{(1)}=Y_{(0)}^{1/2}(Y_{(0)}^{-1/2}Y_{(1)}Y_{(0)}^{-1/2})^{\rho/5}Y_{(0)}^{1/2}.$
  • $\mathbb{S}^{2}$$\mathcal{S}_{2}^{++}$-Hybrid: Here we consider $X\in\mathbb{S}^{2}$ as on the unit sphere, while $Y\in\mathcal{S}_{2}^{++}$ is a symmetric positive definite matrix. First, generate $X_{(0)}$ and $Y$ as in the setting $\mathcal{S}_{2}^{++}$-Interpolation. Then, $X$ is the vectorized form of the upper triangular part of $X_{(0)}$, normalized so that $\|X\|_{2}=1$.
@@
  • $\mathbb{W}^{2}$-Mean: $X$ and $Y$ are random distributions in $\mathbb{W}^{2}$, specifically $X=\mathcal{N}(\mu_{X},\sigma_{X})$ and $Y=\mathcal{N}(\mu_{Y},\sigma_{Y})$, where $\mu_{X}\sim\mathcal{N}(0,1)$, $\mu_{Y}=0.2\epsilon\times|\mu_{X}|^{\rho}$, $\sigma_{X}\sim\mathrm{Unif}(0,1)$, $\sigma_{Y}\sim\mathrm{Unif}(1,2)$ and $\epsilon\sim\mathcal{N}(0,1)$.
We compare the proposed profile-based independence test with the energy test (Székely et al., 2007) and the ball divergence test (Pan et al., 2020). All methods are implemented using available R packages: the energy test via dcov.test() from the energy package, the ball divergence test via bcov.test() from the Ball package (Zhu et al., 2021), and Chatterjee’s rank test via xicor() through the XICOR package. Since Chatterjee’s rank test only supports scalar independence testing, we include it only for the $\mathbb{R}^{1}$ setting. Figure 1 presents power curves as functions of the dependency parameter $\rho$ for the simulation settings described above. The simulation results show that the proposed profile-based test effectively adapts to complex dependency structures where traditional methods tend to fail. For instance, in the Euclidean space $\mathbb{R}^{2}$, when $Y$ and $X$ exhibit linear dependence, the energy test slightly outperforms the profile-based test. However, the proposed method performs better when the dependency becomes logarithmic. In the more complex setting $\mathbb{R}^{1}$-Cir, both the energy test and Chatterjee test fail, while the proposed test achieves good power and outperforms the ball divergence test. Moreover, in more challenging scenarios involving non-Euclidean spaces $\mathcal{S}_{2}^{++}$, $\mathbb{S}^{2}$, and $\mathbb{W}^{2}$, the proposed test performs significantly better than all other comparable methods. These results are consistent with our theoretical findings and highlight the robustness and efficiency of the proposed test across a wide range of dependency structures. In summary, the proposed method performs comparably to existing approaches under linear dependence and outperforms them when the dependency structure is more intricate.
@@
Figure 1: Power curves comparing the proposed profile-based independence test (blue solid lines), the ball divergence test (black dashed lines), the energy test (red dotted lines) and the Chatterjee test (yellow dash-dotted lines), where power is plotted as a function of the dependency parameter $\rho$, defined in the second paragraph of section 5, across six simulation settings: $\mathbb{R}^{1}$-Cir, $\mathbb{R}^{2}$-Log, and $\mathbb{R}^{10}$-Log (left, middle, and right panels in the first row); $\mathcal{S}_{2}^{++}$-Interpolation, $\mathbb{S}^{2}$$\mathcal{S}_{2}^{++}$-Hybrid, $\mathbb{W}^{2}$-Mean (left, middle, and right panels in the second row). The dashed grey line indicates the target significance level of 0.05, and the sample size is set to $n=200$.
We next study the performance of the proposed method in several high-dimensional settings. Figure 2 presents the rejection rates as a function of sample size for the settings $\mathbb{R}^{100}$-Log and $\mathbb{W}^{2}$-Mean under both $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$. The results demonstrate that the proposed method consistently controls the type I error across all sample sizes. In the Euclidean $\mathbb{R}^{100}$-Log setting, the proposed method outperforms the energy test and performs comparably to the ball divergence test, with slightly better power for large sample sizes. This suggests that, although the proposed method is not specifically tailored for high-dimensional data, it still achieves competitive or superior finite-sample performance. In the infinite-dimensional $\mathbb{W}^{2}$-Mean setting, the proposed method exhibits a near-perfect power curve for relatively large sample sizes, while all other methods fail and show limited rejection rates.
We include results for additional settings in the Supplement, illustrating how the power curves evolve with increasing sample size and approach perfect power for large samples. Additionally, we assess the performance of the proposed test across a range of significance levels in the Supplement.
Figure 2: Rejection rate of the proposed independence test for two high-dimensional settings (left two panels: $\mathbb{R}^{100}$-Log, right two panels: $\mathbb{W}^{2}$-Mean). Different colors and linetypes represent different methods: blue solid for the proposed method, black dashed for the ball divergence test, and red dotted for the energy test. Within each setting, two separate plots show the results under $\mathcal{H}_{0}$ (left) and $\mathcal{H}_{1}$ (right). The dashed grey line indicates the target significance level of 0.05
Figure 3: Power curves of the proposed conditional independence test across four simulation settings with different metric spaces: Conditional-$\mathbb{S}^{2}$-Log, Conditional-$\mathbb{S}^{4}$-Log, Conditional-$\mathbb{W}^{2}$-Log, and Conditional-$\mathbb{W}^{2}$-Sin. Each plot shows the rejection rate as a function of the dependency parameter $\rho$, defined in the second paragraph of section 5, with sample size fixed at $n=200$. The covariate $Z$ is drawn from $\mathrm{Unif}[0,1]$, and each point is based on 500 Monte Carlo replications. The dashed grey line indicates the target significance level of 0.05.
For the conditional independence testing, since there are no existing methods designed specifically for random objects, we focus on evaluating the performance of the proposed method in several metric space settings. In all simulation settings below, the covariate $Z$ is generated from a $\mathrm{Unif}[0,1]$ distribution, and $\rho$ denotes the strength of dependence as defined in the second paragraph of section 5,
  • Conditional-$\mathbb{S}^{p}$-Log: $\bm{X}=(X_{1},X_{2},\ldots,X_{p+1})$ and $\bm{Y}=(Y_{1},Y_{2},\ldots,Y_{p+1})$ are random elements on the unit sphere $\mathbb{S}^{p}$, generated with $Z$ and independent standard normal random variable $\epsilon_{1},\,\epsilon_{2},\ldots,\epsilon_{2p+1}$:
    $X_{1}=\frac{Z+\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\ldots,\,X_{p+1}=\frac{\epsilon_{p+1}}{\|\bm{\epsilon}_{X}\|};$
    $Y_{1}=\frac{Z+\rho\log(4\epsilon_{1}^{2})+(1-\rho)\epsilon_{p+2}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{p+3}}{\|\bm{\epsilon}_{Y}\|},\,\ldots,\,\,Y_{3}=\frac{\epsilon_{2p+2}}{\|\bm{\epsilon}_{Y}\|},$
    where $\bm{\epsilon}_{X}=(Z+\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{p+1})^{\top}$, $\bm{\epsilon}_{Y}=(Z+\rho\log(4\epsilon_{1}^{2})+(1-\rho)\epsilon_{p+2},\epsilon_{p+3},\ldots,\epsilon_{2p+2})^{\top}$.
  • Conditional-$\mathbb{W}^{2}$-Log: $X$ and $Y$ are random elements in the space of distributions $\mathbb{W}^{2}$ with the Wasserstein metric, specifically $X=\mathcal{N}(\mu_{X},1)$ and $Y=\mathcal{N}(\mu_{Y},1)$, where $\mu_{X}\sim Z+\mathrm{Unif}(-1,1)$, $\mu_{Y}=Z+\rho\log(4\mu_{X}^{2})+\epsilon$, and $\epsilon\sim\mathrm{Unif}(-1,1)$ and is independent of $\mu_{X}$ and $Z$.
  • Conditional-$\mathbb{W}^{2}$-Sin: $X$ and $Y$ are random elements in the space of distributions $\mathbb{W}^{2}$ with the Wasserstein metric, specifically $X=\mathcal{N}(\mu_{X},1)$ and $Y=\mathcal{N}(\mu_{Y},1)$, where $\mu_{X}\sim Z+\mathrm{Unif}(-1,1)$, $\mu_{Y}=Z+\rho\sin(\mu_{X}\pi)+\epsilon$, and $\epsilon\sim\mathrm{Unif}(-1,1)$ and is independent of $\mu_{X}$ and $Z$.
The results are shown in Figure 3. The proposed conditional test shows increasing power as the dependency parameter $\rho$ grows, achieving near-perfect rejection rates for moderate to strong dependence. These results highlight the adaptability and robustness of the proposed method across different metric spaces and across complex nonlinear scenarios such as the log and sinusoidal dependence structures.